How many 5 letter-words using letters from TRIANGLE can be made if the five-letter word must have two vowels and three consonants?
I set up the question using 5 blank spaces and wrote it like this:
5 4 3 3 2
then I multiplied
and I got 360 different 5 letter words.
answer key says 3600, where did I go wrong?

Answer by Edwin McCravy(20054)


TRIANGLE
has
consonants T,R,N,G,L
vowels I,A,E

We can choose three consonants in C(5,3) ways.
We can choose two vowels in C(3,2) ways.
We can permute the 5 letters in 5! ways:

C(5,3)*C(3,2)*5! = 10*3*120 = 3600

Edwin